OpenStudy (anonymous):

Radius of curvature=35m, linear speed=10km/hr, find the angular velocity and acceleration. Please help me.

OpenStudy (mani_jha):

Use \[v=wR\] Put v=10*5/18 m/s R=35

OpenStudy (anonymous):

Can I use linear Speed=10km/hr=2.778m/s directly substitute in the formula a=v^2/R a=2.778^2/35=0.2208m/s^2. Is this correct?

OpenStudy (vincent-lyon.fr):

Yes it is!

OpenStudy (anonymous):

\[\alpha=a/R \] alpha is the angular acceleration

OpenStudy (vincent-lyon.fr):

@Natan In this example angular acceleration is zero. and (linear) acceleration is purely centripetal. If acceleration had a component tangent to the circle, then we would have: \(\large\alpha=\huge\frac{a_t}{R}\)

OpenStudy (anonymous):

oh sorry !! is true it is a uniform circular motion thanks!